# Unveiling quantum phase transitions from traps in variational quantum algorithms

### v1

@@ Appendix A Exemplary Set Transformer Code
Figure 8: Exemplary code for the core architecture of the Set Transformer, utilized for detecting quantum phase transitions in the eSSH model. This includes implementations of both the Self-Attention Block (SAB) and the overall regression model structure.
Appendix B Robustness of Hyperparameter Sensitivity
@@
This section evaluates the stability and robustness of our hybrid quantum optimization-machine learning algorithm against variations in two key hyperparameters: the regularization parameter $\lambda$ within the LASSO algorithm, which inversely influences the width of the loss function landscape’s valleys, and the window size $w$, crucial for both the LASSO and Transformer-based algorithms for setting the range of phase labels $-1$ and $1$. It is essential to avoid excessively high values of $\lambda$, as this would result in all coefficients $\bm{\kappa}$ shrinking to zero, thereby reducing the LASSO’s target cost function to a constant value of 1. We analyze the algorithm’s performance using three illustrative examples: LASSO for the numerical data from the 1D TFIM, LASSO for the experimental data from the 2D TFIM, and the Set Transformer for the eSSH model with $\delta=4$, where two phase transitions are observed.
@@
The results, displayed in Fig. 9, include Panels (a) and (b) for the 1D TFIM, Panels (c) and (d) for the 2D TFIM, and Panel (e) for the eSSH model. Panel (a) shows the predicted critical points for various valley widths with a constant window size of $w=30$, demonstrating small sensitivity to this hyperparameter in LASSO as all lines remain constant. Panel (b) illustrates the predicted critical points across different window sizes with a set valley width of 0.06, where the maximum deviation observed is only $0.01$. Panel (c) displays results for different valley widths at a window size of $w=4$, consistently converging to $1.6$. Panel (d) examines various window sizes at a valley width of $5$, where the maximum variation is $0.2$, but consistently approximates to the most likely value of $1.6$. Panel (e) evaluates the Transformer’s efficacy on the eSSH model with $\delta=4$, considering various window sizes. It shows that the standard deviations associated with predictions of phase transitions from trivial to symmetry-broken (SB) phases and from SB to topological phases are 0.00525 and 0.00482, respectively, both indicating low variability.
@@
Figure 9: Predicted critical points for varying valley widths and window sizes, demonstrating the algorithm’s robustness with respect to hyperparameters $\lambda$ and $w$ across diverse settings. (a) and (b) focus on processing numerical data for the 1D transverse-field Ising model (TFIM) using LASSO. In both panels, colours correspond to the different values of $p$ defined in Fig. 5(a, b). (c) and (d) present results from processing experimental data for the 2D TFIM using LASSO. (e) showcases the Set Transformer algorithm applied to the extended Su-Schrieffer-Heeger (eSSH) model with $\delta=4$, highlighting its response to changes in $w$.
Appendix C Robustness to Quasi-Global Depolarizing Noise
Proof of Theorem 1  The feature vector $\bm{f}(g)$, input into the machine learning model, consists of the expectation values of Pauli operators from the set $\mathcal{P}_{f}$:
$\bm{f}(g)=\left(\operatorname{Tr}\left(O_{1}\rho(\bm{\zeta},\bm{\eta};g)\right),\ldots,\operatorname{Tr}\left(O_{\ell}\rho(\bm{\zeta},\bm{\eta};g)\right)\right),$ (36)
@@ Appendix C Robustness to Quasi-Global Depolarizing Noise
$\Lambda_{\epsilon}(\rho)=(1-\epsilon)\rho+\epsilon\frac{I}{2^{n}},$ (37)
@@
$\operatorname{Tr}\left(O\Lambda_{\epsilon}(\rho(\bm{\zeta},\bm{\eta};g))\right)=(1-\epsilon)\operatorname{Tr}(O\rho(\bm{\zeta},\bm{\eta};g))+\frac{\epsilon}{2^{n}}\operatorname{Tr}\left(O\right).$ (38)
@@
$\operatorname{Tr}\left(O\Lambda_{\epsilon}(\rho(\bm{\zeta},\bm{\eta};g))\right)=(1-\epsilon)\operatorname{Tr}(O\rho(\bm{\zeta},\bm{\eta};g)).$ (39)
Consequently, the noisy feature vector $\bm{f}_{\text{noisy}}(g)$ is a scaled version of the original feature vector:
$\bm{f}_{\text{noisy}}(g)=(1-\epsilon)\bm{f}(g).$ (40)
When the noisy feature vector $\bm{f}_{\text{noisy}}(g)$ is processed using LASSO, we can adjust the regularization coefficient $\lambda$ by multiplying it by $(1-\epsilon)$. Considering the LASSO cost function as shown in Eq. (1):
$\mathcal{C}(\bm{\kappa},\lambda)=\left(\frac{1}{4w}\sum_{i=1}^{2w}\left(l_{i}-\kappa_{0}-\sum_{j=1}^{\ell}\kappa_{j}f_{ij}\right)^{2}+\lambda\sum_{j=1}^{\ell}\left|\kappa_{j}\right|\right)$ (41)
for the coefficients $\bm{\kappa}=\left(\kappa_{0},\kappa_{1},\kappa_{2},\dots,\kappa_{\ell}\right)$, the same cost function $\mathcal{C}$ can be achieved with the input $\bm{f}(g)$ and the adjusted coefficients:
$\bm{\kappa}_{\text{noisy}}=\left(\kappa_{0},\frac{\kappa_{1}}{1-\epsilon},\frac{\kappa_{2}}{1-\epsilon},\dots,\frac{\kappa_{\ell}}{1-\epsilon}\right).$ (42)
@@
When processing the noisy feature vector $\bm{f}_{\text{noisy}}(g)$ through the Transformer model, the initial step involves normalization of feature vectors. This step rescales data and effectively mitigating the scaling introduced by the noise, ensuring that the learning model’s output remains unaffected by the scaling effect of the quasi-global depolarizing noise. $\blacksquare$

### v2

@@ VI Larger System Investigations for the 2D TFIM
Figure S3: Exemplary code for the core architecture of the Set Transformer, utilized for detecting quantum phase transitions in the eSSH model. This includes implementations of both the Self-Attention Block (SAB) and the overall regression model structure.
In panel (a) of Fig. S2, we plot the derivatives of the absolute relative spectral gap $(E_{1}-E_{\text{gs}})/|E_{\text{gs}}|$ as a function of the transverse field $g$ for the aforementioned system sizes. The estimated critical points from these derivatives are approximately 1.13, 1.96, 2.88, 2.96, 3.04, and 3.06 for the $2\times 2$, $2\times 3$, $3\times 3$, $3\times 4$, $4\times 4$, and $4\times 5$ systems, respectively. Notably, the last two estimates are very close to the thermodynamic limit value, $g_{c}=3.04438$ [71, 72]. Panel (b) shows the derivatives of the $x$-magnetization as a function of $g$ for the same set of system sizes. The corresponding estimated critical points are 0.68, 1.34, 2.15, 2.32, 2.48, and 2.57. These estimates are clearly less accurate than those obtained from the relative spectral gap, highlighting the limitations of local order parameters in finite systems.
We further implement our LASSO-based algorithm on a $4\times 4$ grid with periodic boundary conditions for circuit depths $p=2,3,4$, and $5$. The average fidelities for these depths are 0.5473, 0.7336, 0.8674, and 0.9293, respectively. In panel (c) of Fig. S2, we display the LASSO loss landscapes for each circuit depth. The estimated critical points from the loss landscapes are 1.26, 2.11, 2.38, and 2.45 for $p=2,3,4$, and $5$, respectively. For comparison, we also include, using dashed lines, the critical boundaries obtained from the $x$-magnetization of the trapped states. Notably, for $p=2$, the $x$-magnetization fails to provide an estimate, and for $p=3,4,5$, the estimates derived from $x$-magnetization are considerably worse than those obtained via our algorithm. The dominant order parameter identified by our method in this case is $\mathcal{O}_{\mathrm{ml}}=\sum_{i,j}\sigma_{i,j}^{x}\sigma_{i+2,j+1}^{x}$.
Finally, in panel (d), we present the derivative of the learned order parameter (i.e., $\mathcal{O}_{\mathrm{ml}}$) with respect to $g$, computed from the ground state data for both the $4\times 4$ and $4\times 5$ grids. The derivative curves obtained using the learned order parameter yield more accurate estimates of the critical point (2.56 and 2.62, respectively) compared to those based on $x$-magnetization.
VII Exemplary Set Transformer Code and Complexity Analysis
In Fig. S3 we provide the exemplary Python code for the Set Transformer used in Section II.3. Central to this architecture is the Self-Attention Block (SAB), which processes the input data through the mechanism of self-attention. In this setup, the normalized input, $X\_{\text{norm}}$, serves simultaneously as the Query, Key, and Value. This configuration enables each element of the input feature interact with every other feature, facilitating a comprehensive internal representation that captures the underlying relationships.
Figure S4: Predicted critical points for varying valley widths and window sizes, demonstrating the algorithm’s robustness with respect to hyperparameters $\lambda$ and $w$ across diverse settings. (a) and (b) focus on processing numerical data for the 1D transverse-field Ising model (TFIM) using LASSO. In both panels, colours correspond to the different values of $p$ defined in Fig. 3(a, b). (c) and (d) present results from processing experimental data for the 2D TFIM using LASSO. (e) showcases the Set Transformer algorithm applied to the extended Su-Schrieffer-Heeger (eSSH) model with $\delta=4$, highlighting its response to changes in $w$.
In our proposed framework, a set of optimization grid points, whose total number we denote by $N$, is chosen as potential candidates for detecting phase transitions. For each valid center $g$, the algorithm collects a local dataset of size $2w$, formed by $w$ parameter values below $g$ (labeled $-1$) and $w$ parameter values above $g$ (labeled $+1$), and then trains a supervised model on these samples. Let $\ell=\ell(n)$ be the total number of observables (Pauli operators or otherwise) whose expectation values are measured and fed into the model. In practice, $\ell$ may be chosen to grow polynomially in the system size $n$ if only local or few-body operators are included, or it may grow exponentially when higher-body correlations are considered.
In each training session, we fit a classical regression model—LASSO or a Transformer—to a labeled dataset of size $2w$. For LASSO, we employ a coordinate descent solver, which incurs a per-iteration computational cost of order $w\ell^{2}$. If the algorithm requires $s$ iterations to converge, the cost of a single LASSO training run is $\mathcal{O}(s\,w\,\ell^{2})$. When $s$ and $w$ remain nearly constant or grow only slowly with $n$, $\ell(n)^{2}$ becomes the dominating factor.
For Transformers inspired by the Set Transformer framework, the forward/backward pass of a single attention layer scales as $\mathcal{O}(\ell^{2}\,d)$, where $d$ is the embedding dimension. In the presence of $h$ attention heads, the overall cost for one layer is approximately $\mathcal{O}(h\,\ell^{2}\,d)$. If $\tau$ is the number of epochs or gradient updates, each Transformer training session has an overall cost of $\mathcal{O}(\tau\,h\,\ell^{2}\,d)$, possibly with additional constants from feedforward blocks. Because the dataset size is $2w$, $\tau$ generally stays modest in most practical settings, so again $\ell(n)^{2}$ often drives the scaling.
Since the sliding-window procedure trains a separate model at each center $g$, these costs must be multiplied by the number of centers $N_{c}$. If boundary effects in the parameter grid are negligible, $N_{c}$ is of the same order as $N$. Consequently, applying LASSO at all window centers yields a total cost of $\mathcal{O}\bigl{(}N_{c}\times s\,w\,\ell^{2}\bigr{)}$, while a Transformer-based approach costs $\mathcal{O}\bigl{(}N_{c}\times\tau\,h\,\ell^{2}\,d\bigr{)}$. In typical setups, $s$, $\tau$, $h$, $d$, and $w$ are constants or grow only mildly with $n$, making $\ell(n)^{2}$ and $N_{c}$ the primary contributors to the runtime. If $\ell(n)$ can be chosen to scale polynomially with $n$ (or remain approximately constant), and $N_{c}$ is a constant independent of $n$, then the overall complexity of the sliding-window detection algorithm scales polynomially with $n$. Under these assumptions, the protocol remains computationally feasible for moderate system sizes relevant to current or near-term quantum simulations.
Finally, we empirically evaluated the runtime of our LASSO-based detection procedure on the 1D TFIM for system sizes $n=4,6,\ldots,16$. As shown in Table S1, we fixed the window size at $w=30$ and the number of sampled observables at $\ell=30$, and scanned $N_{c}=1940$ boundary points. Despite increasing $n$, the total runtime remains nearly constant at around 1–1.4 seconds, indicating that once $w$ and $\ell$ are fixed, our post-processing overhead does not appreciably grow with the system size.
System size 4 6 8 10 12 14 16
Runtime (s) 1.381 1.395 1.308 1.405 1.381 1.296 1.405
Table S1: Total runtime (in seconds) for the LASSO-based detection of phase boundaries in the 1D TFIM, with $N_{c}=1940$, $\ell=30$, and $w=30$ kept fixed, across various system sizes $n$.
From a theoretical standpoint, the iteration count or epoch number in these “black box” ML models depends largely on the complexity of distinguishing phases in feature space rather than simply on the system size. In many physically relevant models, this complexity does not necessarily scale with $n$. Consequently, even in regimes where a full classical simulation of the quantum state becomes intractable, the ML-based post-processing is expected remain tractable, provided it only requires distinguishing phases based on a polynomial number of measured features.
VIII Robustness of Hyperparameter Sensitivity
@@
This section evaluates the stability and robustness of our hybrid quantum optimization-machine learning algorithm against variations in two key hyperparameters: the regularization parameter $\lambda$ within the LASSO algorithm, which inversely influences the width of the loss function landscape’s valleys, and the window size $w$, crucial for both the LASSO and Transformer-based algorithms for setting the range of phase labels $-1$ and $1$. It is essential to avoid excessively high values of $\lambda$, as this would result in all coefficients $\boldsymbol{\kappa}$ shrinking to zero, thereby reducing the LASSO’s target cost function to a constant value of 1. We analyse the algorithm’s performance using three illustrative examples: LASSO for the numerical data from the 1D TFIM, LASSO for the experimental data from the 2D TFIM, and the Set Transformer for the eSSH model with $\delta=4$, where two phase transitions are observed.
@@
The results, displayed in Fig. S4, include panels (a) and (b) for the 1D TFIM, panels (c) and (d) for the 2D TFIM, and panel (e) for the eSSH model. Panel (a) shows the predicted critical points for various valley widths with a constant window size of $w=30$, demonstrating small sensitivity to this hyperparameter in LASSO as all lines remain constant. Panel (b) illustrates the predicted critical points across different window sizes with a set valley width of 0.06, where the maximum deviation observed is only $0.01$. Panel (c) displays results for different valley widths at a window size of $w=4$, consistently converging to $1.6$. Panel (d) examines various window sizes at a valley width of $5$, where the maximum variation is $0.2$, but consistently approximates to the most likely value of $1.6$. Panel (e) evaluates the Transformer’s efficacy on the eSSH model with $\delta=4$, considering various window sizes. It shows that the standard deviations associated with predictions of phase transitions from trivial to symmetry-broken (SB) phases and from SB to topological phases are 0.00525 and 0.00482, respectively, both indicating low variability.
@@
Figure S5: Effect of local depolarizing noise on Pauli expectation ratios in Hamiltonian variational ansatz circuits for the 1D TFIM. The four panels correspond to a different circuit depth: (a) $p=1$, (b) $p=2$, (c) $p=3$, and (d) $p=4$. For each panel, the horizontal axis represents the local depolarizing noise rate $\epsilon^{(\text{DP})}$ and the vertical axis displays the ratio $R(O_{j},\rho)=\operatorname{Tr}\left(O_{j}\mathcal{N}_{\text{DP}}(\rho)\right)/\operatorname{Tr}\left(O_{j}\rho\right)$ for a randomly chosen Pauli operator $O_{j}$. Each data point is associated to the state $\rho$ corresponding to one of 200 quantum circuits with a randomly sampled set of rotation angles $\{\boldsymbol{\beta},\boldsymbol{\gamma}\}$. Blue points correspond to measurements of weight-1 Pauli operators, while red points correspond to weight-2 Pauli operators. The results demonstrate that for shallower circuits and lower noise rates, the variance of the ratio is minimal, validating the quasi-global depolarizing noise approximation for the randomly chosen observables.
IX Robustness to Quasi-Global Depolarizing Noise
Proof of Theorem 1  The feature vector $\boldsymbol{f}(g)$, input into the machine learning model, consists of the expectation values of Pauli operators from the set $\mathcal{P}_{f}$:
$\boldsymbol{f}(g)=\left(\operatorname{Tr}\left(O_{1}\rho(\boldsymbol{\zeta},\boldsymbol{\eta};g)\right),\ldots,\operatorname{Tr}\left(O_{\ell}\rho(\boldsymbol{\zeta},\boldsymbol{\eta};g)\right)\right),$ (S3)
@@ IX Robustness to Quasi-Global Depolarizing Noise
$\Lambda_{\epsilon}(\rho)=(1-\epsilon)\rho+\epsilon\frac{I}{2^{n}},$ (S4)
@@
$\operatorname{Tr}\left(O\Lambda_{\epsilon}(\rho(\boldsymbol{\zeta},\boldsymbol{\eta};g))\right)=(1-\epsilon)\operatorname{Tr}(O\rho(\boldsymbol{\zeta},\boldsymbol{\eta};g))+\frac{\epsilon}{2^{n}}\operatorname{Tr}\left(O\right).$ (S5)
@@
$\operatorname{Tr}\left(O\Lambda_{\epsilon}(\rho(\boldsymbol{\zeta},\boldsymbol{\eta};g))\right)=(1-\epsilon)\operatorname{Tr}(O\rho(\boldsymbol{\zeta},\boldsymbol{\eta};g)).$ (S6)
Consequently, the noisy feature vector $\boldsymbol{f}_{\text{noisy}}(g)$ is a scaled version of the original feature vector:
$\boldsymbol{f}_{\text{noisy}}(g)=(1-\epsilon)\boldsymbol{f}(g).$ (S7)
When the noisy feature vector $\boldsymbol{f}_{\text{noisy}}(g)$ is processed using LASSO, we can adjust the regularization coefficient $\lambda$ by multiplying it by $(1-\epsilon)$. Considering the LASSO cost function as shown in Eq. (26):
$\mathcal{C}(\boldsymbol{\kappa},\lambda)=\left(\frac{1}{4w}\sum_{i=1}^{2w}\left(l_{i}-\kappa_{0}-\sum_{j=1}^{\ell}\kappa_{j}f_{ij}\right)^{2}+\lambda\sum_{j=1}^{\ell}\left|\kappa_{j}\right|\right)$ (S8)
for the coefficients $\boldsymbol{\kappa}=\left(\kappa_{0},\kappa_{1},\kappa_{2},\dots,\kappa_{\ell}\right)$, the same cost function $\mathcal{C}$ can be achieved with the input $\boldsymbol{f}(g)$ and the adjusted coefficients:
$\boldsymbol{\kappa}_{\text{noisy}}=\left(\kappa_{0},\frac{\kappa_{1}}{1-\epsilon},\frac{\kappa_{2}}{1-\epsilon},\dots,\frac{\kappa_{\ell}}{1-\epsilon}\right).$ (S9)
@@
When processing the noisy feature vector $\boldsymbol{f}_{\text{noisy}}(g)$ through the Transformer model, the initial step involves normalization of feature vectors. This step rescales data and effectively mitigating the scaling introduced by the noise, ensuring that the learning model’s output remains unaffected by the scaling effect of the quasi-global depolarizing noise. $\blacksquare$
To empirically validate the conditions specified in Theorem 1, we conducted numerical simulations using the Hamiltonian variational ansatz designed for the one-dimensional transverse-field Ising model (1D TFIM). In our simulation framework, each two-qubit rotation gate is subject to local depolarizing noise with a rate of $\epsilon^{(\text{DP})}$, which results in an effective two-qubit error rate of approximately $2\epsilon^{(\text{DP})}$.
We generated a representative ensemble of circuit outputs by randomly sampling 200 distinct sets of rotation angles $\{\boldsymbol{\beta},\boldsymbol{\gamma}\}$. For each resultant quantum state $\rho$, we randomly select one Pauli operator $O_{j}$ from the set of Pauli operators with weight 1 or 2. To ensure that only significant contributions are considered, we imposed the condition $|\operatorname{Tr}(O_{j}\rho)|>0.2$. This thresholding procedure serves to filter out Pauli terms with negligible expectation values, which would otherwise be disproportionately affected by small noise-induced fluctuations.
The key metric we analyse is the ratio of the expectation value of the noisy state to that of the ideal (noiseless) state, defined as $R(O_{j},\rho)=\operatorname{Tr}\left(O_{j}\mathcal{N}(\rho)\right)/\operatorname{Tr}\left(O_{j}\rho\right)$. We evaluated this ratio for a system size of $n=10$ qubits and varying circuit depths $p=1,2,3,4$, while systematically adjusting the gate noise rate to different values $\epsilon^{(\text{DP})}=0.0025,0.005,0.0075,0.01,0.0125,0.015$. The results, depicted in Fig. S5, display the distribution of the computed ratios for different configurations. Each subplot corresponds to a specific circuit depth, with blue markers representing ratios for weight-1 Pauli operators and red markers representing ratios for weight-2 Pauli operators.
We observe that the variance of the computed ratios is smaller when: (i) the circuit depth is low, (ii) the noise rate is small, and (iii) the observables are low-weight. The third behavior can be understood theoretically by noting that, for shallow circuits, the noise affecting individual qubits can be treated as statistically independent. For an operator that acts nontrivially on $k$ qubits, the overall noise-induced attenuation is approximately given by the product $(1-\epsilon^{\prime})^{k}$ for certain $\epsilon^{\prime}$. For low-weight operators (small $k$), the expectation values are only marginally reduced and exhibit low variance. In contrast, for high-weight operators, the multiplicative noise effects compound, leading to significantly lower expectation values and higher relative variance. As the circuit depth increases, the variance in the ratios grows, indicating that the quasi-global depolarizing approximation underlying Theorem 1 becomes progressively less valid. These results imply that by carefully selecting circuit depths and favoring lower-weight Pauli operators for $\mathcal{P}_{f}$, the impact of noise can be more effectively mitigated in practical implementations.